A parallel plate capacitor having a plate area of 0.70 m2 and a plate separation of 1.0 mm is connected to a source with a voltage of 50 V. Find the capacitance, the charge on the plates, and the energy of the capacitor b) when the capacitor contains polystyrene


1
Expert's answer
2021-03-10T17:15:54-0500

1. The capacitance of the parallel plate capacitor is given as follows:


"C = \dfrac{\varepsilon_0\varepsilon S}{d}"

where "\varepsilon_0 = 8.85 \cdot 10^{−12}F/m" is the electric constant, "\varepsilon = 2.5" is the permittivity of polystyrene, "S = 0.7m^2" is the area of the plates, and "d = 1mm = 10^{-3}m" is the distance between the plates.

Thus, obtain:


"C = \dfrac{8.85 \cdot 10^{−12}\cdot 2.5\cdot 0.7}{10^{-3}} \approx 15.5\times10^{-9}F"

2. The charge on the plates (on one plate) is:


"q = CV"

where "V = 50V" is the voltage of the source. Thus, obtain:


"q = 15.5\times 10^{-9}F\cdot 50V \approx 7.8\times 10^{-7}C"

3. The energy stored in the capacitors is:


"W = \dfrac{CV^2}{2}\\
W = \dfrac{15.5\times10^{-9}\times 50^2}{2}\approx 1.9\times 10^{-5}J"

Answer. a) "15.5\times10^{-9}F", b) "7.8\times 10^{-7}C", c) "1.9\times 10^{-5}J".